# From Open Vocabulary to Open World: Teaching Vision Language Models to Detect Novel Objects

### v3

@@ Abstract
Traditional object detection methods operate under the closed-set assumption, where models can only detect a fixed number of objects predefined in the training set. Recent works on open vocabulary object detection (OVD) enable the detection of objects defined by an in-principle unbounded vocabulary, which reduces the cost of training models for specific tasks. However, OVD heavily relies on accurate prompts provided by an “oracle”, which limits their use in critical applications such as driving scene perception. OVD models tend to misclassify near-out-of-distribution (NOOD) objects that have similar features to known classes, and ignore far-out-of-distribution (FOOD) objects. To address these limitations, we propose a framework that enables OVD models to operate in open world settings, by identifying and incrementally learning previously unseen objects. To detect FOOD objects, we propose Open World Embedding Learning (OWEL) and introduce the concept of Pseudo Unknown Embedding which infers the location of unknown classes in a continuous semantic space based on the information of known classes. We also propose Multi-Scale Contrastive Anchor Learning (MSCAL), which enables the identification of misclassified unknown objects by promoting the intra-class consistency of object embeddings at different scales. The proposed method achieves state-of-the-art performance on standard open world object detection and autonomous driving benchmarks while maintaining its open vocabulary object detection capability.
@@ 1 Introduction
Object detection is a fundamental computer vision task, which involves localization and classification of foreground objects. Although there has been significant progress in this area, many methods rely on the closed-set assumption [12, 50, 11, 49, 30, 28, 61, 3], where all object categories to be predicted are available in the training set.
@@
In many real world applications, such as autonomous driving, the closed-set assumption is unrealistic and even dangerous, because it forces the model to misclassify or ignore unknown objects [6]. Scheirer et al. [53] defines the problem of rejecting unknown classes ($\mathcal{U}$) and simultaneously classifying known classes ($\mathcal{K}$) as open set recognition (OSR). Subsequent works [41, 6] extend this problem to open set object detection (OSOD). Joseph et al. [19] further proposes Open World Object Detection (OWOD), which involves detecting both known and unknown objects and incrementally learning new classes.
Figure 1: Yolo World fails to correctly detect objects that are not included in the prompt. When we use a prompt set comprising all PASCAL VOC classes, the model misclassifies near-out-of-distribution objects (teddy bears) and ignores far-out-of-distribution objects (plates).
@@
Open world object detection is a challenging task due to the complexity of both open-set recognition [41, 6] and incremental learning [57]. The model must generalize beyond predefined classes to capture the objectness of diverse unknown objects, and avoid confusing them with known classes. Additionally, it needs to incrementally learn new classes without forgetting previously acquired knowledge. Despite some progress [19, 14, 69, 68, 36, 75, 34, 65, 59] in this area, several key issues still remain unresolved. Many existing methods perform poorly at discovering unknown objects, leading to low recall for unknown classes. Additionally, existing OWOD methods [19, 14, 59, 65, 75, 36, 69, 68, 34] employ a replay strategy during incremental learning, where data from previous tasks are reintroduced during the training of new classes. This is not an optimal strategy as it wastes computing power and time.
@@
Recently, the rise of visual-language pre-training has spawned a new area: open vocabulary object detection (OVD). In principle, OVD can detect novel classes defined by an unbounded (open) vocabulary at inference. Naturally, it is able to incrementally learn new classes by adding new prompts. In the real world, however, the open vocabulary may not necessarily be complete and accurate. As shown in Fig. 1, an open vocabulary model such as Yolo World [5] is still a closed-set model at test time, because it is impractical for the expert user, referred to as “oracle”, to define all objects that the model might encounter in the real world. With a limited vocabulary the model tends to misclassify near-out-of-distribution (NOOD) objects that have similar semantics to known classes, and ignore far-out-of-distribution (FOOD) objects. In real-world applications such as autonomous driving, where countless object-types cannot be initially included in the text prompt, OVD will inevitably fail when presented with OOD objects.
@@
To identify NOOD objects, we propose Multi-Scale Contrastive Anchor Learning (MSCAL). Contrastive learning has been used in OOD detection [60, 43, 33] to pull similar samples together and push dissimilar samples apart in the representation space. The key insight of MSCAL is that in an open world setting, object classes are introduced gradually. As new classes are introduced, the decision boundaries for known classes shift within the shared feature space. We formulate the task of unknown object identification in the open-set setting as a series of deep one-class classification [51] problems. For each class $i$, we use an individual non-linear projector to map the feature pyramid into a class-specific representation space, contrasting embeddings with class anchors. Positive samples from class $i$, at different scales, maximize their similarity to the class anchor, while embeddings from other classes and the background act as negative samples. MSCAL ensures that embeddings of known classes at different scales are tightly clustered around their corresponding anchors, while unknown object embeddings are left out of the clusters and can be rejected by a distance based OOD score.
Our method significantly surpasses state-of-the-art (SOTA) performance in U-Recall on the M-OWODB [19] and S-OWODB [14] benchmarks, while maintaining leading performance in other metrics. We further evaluate our method on a novel benchmark based on nuScenes [2], where it also achieves the best results. More importantly, our method preserves the zero-shot capability of the OVD model, as our implementation only optimizes text embeddings and additional MSCAL modules, while keeping the OVD model’s weights frozen. Therefore, our method provides a unified framework for both open vocabulary learning and open world learning. The contributions of this work are as follows:
@@ 2 Related Works
Closed set object detection has been extensively studied over the past decade [12, 50, 11, 49, 30, 28, 61, 3]. To handle unknown objects and incrementally learn new classes, Joseph et al. [19] first proposed an open world object detector, which extends the Faster R-CNN [50] with contrastive clustering and energy-based unknown-object identification. Subsequent works [14, 36] used contextual information to improve unknown identification and knowledge transfer between known and unknown classes. To effectively detect unknown objects, Zohar et al. [75] proposed Probabilistic Objectness (PROB) to estimate the objectness of different proposals. Wang et al. [65] introduced random proposals in detector training to encourage the unknown discovery and reduce confusion between known and unknown classes. Ma et al. [34] proposed decoupling object localization and classification via cascade decoding. Sun et al. [59] further de-correlates objectness and class information by enforcing orthogonality. In some recent works [35, 37, 74], foundation models are employed in open-world learning.
Open Vocabulary Object Detection
@@
Open vocabulary object detection (OVD) aims to detect novel classes with the help of vocabulary knowledge [66]. OVD methods are implemented in various ways, including knowledge distillation [13, 38, 47], region-text pretraining [23, 71, 5] and prompt modeling [67, 7]. An OVD model matches the image embeddings with the text embeddings. When presented with an object that is not in the prompt, the model will either assign an incorrect label of the nearest match (NOOD case) or no detection at all (FOOD case).
OOD Detection with Vision Language Model
Out-of-distribution (OOD) detection is a concept closely related to open set learning, and the two terms are sometimes used interchangeably. OOD detection differs slightly from OSR on the choice of testing data [10, 52].
OOD detection has been extensively studied in standard deep neural networks [42, 16, 31, 62, 43, 26, 22, 58, 60, 76]. Ming et al. [42] introduced the Maximum Concept Matching (MCM) score to assess the OOD-ness of outputs from CLIP [46]. The score can also be extended with negative labels [17]. Capitalizing on the alignment between text and image feature spaces achieved during vision-language pre-training, Wang et al. [63] proposed a prompt learning method for zero-shot OOD detection. Local Regularized Context Optimization (LoCoOp) [44] uses image elements unrelated to in-distribution (ID) samples, such as the background, as negative samples for prompt learning. Similarly, Nie et al. [45] adopted a CoCoOP [72]-style approach to learn negative prompts. Li et al. [24] introduced a transferable negative prompt learning technique tailored to open vocabulary learning, allowing adaptation to new categories.
Contrastive Learning
Contrastive learning aims to learn a representation space where similar instances are close together in the representation space, while dissimilar instances are far apart. Initially, contrastive learning was used to train image encoders in a self-supervised manner [15, 4]. Khosla et al. [21] proposed supervised contrastive learning, which improves the robustness of neural networks. Supervised contrastive learning encourages compactness within each class and separation between different classes, making it suitable for distance-based OOD detection [43]. Some methods [60, 58, 54] iteratively select a sample as an anchor and contrast it with other samples in the mini-batch. More recent approaches [43, 33, 25] employ class anchors, which maximizes the similarity between positive samples and the corresponding class anchor.
@@ 3 Method
Figure 2: Overview of the proposed method. During model training, we first initialize known class embeddings with a pretrained CLIP text encoder [46]. The image encoder extracts a multi-scale feature map from the input. Then the RepVL-PAN [5] uses multi-level cross-modal fusion to combine image and text features, forming the feature pyramids. The detection head predicts the class label based on image-text similarity and regresses the bounding box. The detection loss is used to update the known class embeddings. Concurrently, MSCAL modules are trained to maximize the similarity between class anchor and spatial locations at different scales, and output a multi-scale score map to indicate whether an embedding is out-of-distribution (OOD) relative to a specified class. During the inference, the OOD map extracted by MSCAL is used to reduce known-unknown confusion. In addition, the pseudo unknown embedding used to discover unknown classes is constructed from the optimized known class embeddings and the generic “objectness” semantic concept.
@@
Open World Object Detection (OWOD) [19] aims to detect both known and unknown objects while continuously learning new classes. At an arbitrary stage $t$, we consider the known classes as $\mathcal{K}^{t}=\{1,...,N\}$, and unknown classes as $\mathcal{U}$. An OWOD model should be able to detect objects in $\mathcal{K}^{t}$ and $\mathcal{U}$, and extend known classes to $\mathcal{K}^{t+1}=\mathcal{K}^{t}\cup\{N+1,...,N+k\}$ when $k$ new classes are incrementally learned. In this way, the object detector continuously discovers and learns new classes in the open world.
@@
Fig. 2 shows the general architecture of the proposed method. Following [5], we use text $T$ and image $I$ as inputs and match the text embeddings with image embeddings to predict class labels and bounding boxes of objects. Let $W_{\mathcal{K}}=\{w_{1},...,w_{N}\}$ denote the text embedding of $N$ known classes, which is initialized from class names encoded by the pre-trained CLIP [46] text encoder. $W_{\mathcal{K}}$ can be parameterized as model’s weight and optimized via Open World Embedding Learning (OWEL). During inference, a pseudo unknown embedding is constructed and appended to $W_{\mathcal{K}}$, and the CLIP text encoder is disposable. The image encoder (DarkNet backbone inherited from Yolo v8 [18, 48]) extracts multi-scale features $C$ from the input image $I$. Then the multi-modal neck (RepVL-PAN [5]) uses multi-level cross-modal fusion to combine image and text features, forming the feature pyramids $P$. The detection head predicts bounding boxes and class labels by matching the cosine similarity of text embedding with each spatial location in $P$. Concurrently, MSCAL modules make dense predictions of out-of-distribution (OOD) scores and reject unknown bounding boxes in the detection head. Finally, the redundant predictions are filtered by Non-Maximum Suppression (NMS).
@@
Figure 3: Inferring the Pseudo Unknown Embedding in the embedding space. For CLIP-like models, text embeddings are mapped on a unit hypersphere. The distance between the embeddings reflects the semantic similarity. In a continuous language space, there should be an embedding that represents the generic objectness. Since we know the embeddings of known classes, we can use the generic objectness as a pivot to estimate the Pseudo Unknown Embedding.
@@
Large pre-trained vision-language models have spawned many prompt learning methods [73, 72, 1], which optimize text prompts rather than fine-tune the entire model to improve performance in downstream tasks. But they are not suitable for OWOD, because known $\mathcal{K}$ and unknown $\mathcal{U}$ classes are ever-changing. To address this, we propose a simple and effective way to learn new classes and detect unknown objects, called Open World Embedding Learning (OWEL). For $N$ known classes, we initialize known class embeddings $W_{\mathcal{K}}=\{w_{1},...,w_{N}\}$, and optimize them with the object detection loss. When $k$ new classes are introduced, we simply freeze $W_{\mathcal{K}}$ and train new class embeddings.
@@
As shown in Fig. 1, when an OVD model is given an unknown object, it will either misclassify it as a semantically similar text prompt, or ignore it if the semantic difference is large. To avoid misclassification, we can model the distribution of known classes to reject out-of-distribution samples. To discover novel objects, we construct a Pseudo Class Embedding $w_{\mathcal{U}}$ representing unknown classes.
@@
Vision-language pretraining aligns the visual embedding space and the textual embedding space, such that the text embedding is equivalent to the corresponding image embedding [46, 73, 63, 9]. It has been shown that the relation between text embeddings can be derived through vector offsets in continuous space language models [40, 39]. As shown in Fig. 3, let $\overline{w}$ denote the mean text embedding of all known classes, defined as:
$\overline{w}=\sum_{i=1}^{N}\frac{w_{i}}{N||w_{i}||}.$ (1)
Let $w_{0}$ denote the text embedding representing generic objectness. Naturally, the text prompt of $w_{0}$ should be some general words, such as “object”, which is supported by a previous work [37] and our observation on Yolo World.
@@
The semantics of $w_{0}$ overlaps with known classes to some extent. To shift its focus to unknown classes rather than making duplicate detections on known classes, we propose to construct a Pseudo Unknown Embedding, which is specialized to detect far-out-of-distribution (FOOD) objects. The Pseudo Unknown Embedding $w_{\mathcal{U}}$ is defined by subtracting the mean of known classes from $w_{0}$:
$w_{\mathcal{U}}=w_{0}-\alpha\frac{\overline{w}}{||\overline{w}||},$ (2)

### v4

# Parkville, VIC, Australia From Open Vocabulary to Open World

@@ Abstract
Traditional object detection methods operate under the closed-set assumption, where models can only detect a fixed number of objects predefined in the training set. Recent works on open vocabulary object detection (OVD) enable the detection of objects defined by an in-principle unbounded vocabulary, which reduces the cost of training models for specific tasks. However, OVD heavily relies on accurate prompts provided by an “oracle”, which limits their use in critical applications such as driving scene perception. OVD models tend to misclassify near-out-of-distribution (NOOD) objects that have similar features to known classes, and ignore far-out-of-distribution (FOOD) objects. To address these limitations, we propose a framework that enables OVD models to operate in open world settings, by identifying and incrementally learning previously unseen objects. To detect FOOD objects, we propose Open World Embedding Learning (OWEL) and introduce the concept of Pseudo Unknown Embedding which infers the location of unknown classes in a continuous semantic space based on the information of known classes. We also propose Multi-Scale Contrastive Anchor Learning (MSCAL), which enables the identification of misclassified unknown objects by promoting the intra-class consistency of object embeddings at different scales. The proposed method achieves state-of-the-art performance on standard open world object detection and autonomous driving benchmarks while maintaining its open vocabulary object detection capability. The code is available at https://github.com/343gltysprk/ovow.
@@ 1 Introduction
Object detection is a fundamental computer vision task, which involves localization and classification of foreground objects. Although there has been significant progress in this area, many methods rely on the closed-set assumption [Girshick et al.(2014)Girshick, Donahue, Darrell, and Malik, Ren et al.(2015)Ren, He, Girshick, and Sun, Girshick(2015), Redmon et al.(2016)Redmon, Divvala, Girshick, and Farhadi, Liu et al.(2016)Liu, Anguelov, Erhan, Szegedy, Reed, Fu, and Berg, Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollar, Tian et al.(2019)Tian, Shen, Chen, and He, Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko], where all object categories to be predicted are available in the training set.
@@
In many real world applications, such as autonomous driving, the closed-set assumption is unrealistic and even dangerous, because it forces the model to misclassify or ignore unknown objects [Dhamija et al.(2020)Dhamija, Günther, Ventura, and Boult]. $\bmvaOneDot[Scheirer et al$.(2013)Scheirer, de Rezende Rocha, Sapkota, and Boult] defines the problem of rejecting unknown classes ($\mathcal{U}$) and simultaneously classifying known classes ($\mathcal{K}$) as open set recognition (OSR). Subsequent works [Miller et al.(2018)Miller, Nicholson, Dayoub, and Sünderhauf, Dhamija et al.(2020)Dhamija, Günther, Ventura, and Boult] extend this problem to open set object detection (OSOD). $\bmvaOneDot [Joseph et al$.(2021)Joseph, Khan, Khan, and Balasubramanian] further proposes Open World Object Detection (OWOD), which involves detecting both known and unknown objects and incrementally learning new classes.
@@
Open world object detection is a challenging task due to the complexity of both open-set recognition [Miller et al.(2018)Miller, Nicholson, Dayoub, and Sünderhauf, Dhamija et al.(2020)Dhamija, Günther, Ventura, and Boult] and incremental learning [Shmelkov et al.(2017)Shmelkov, Schmid, and Alahari]. The model must generalize beyond predefined classes to capture the objectness of diverse unknown objects, and avoid confusing them with known classes. Additionally, it needs to incrementally learn new classes without forgetting previously acquired knowledge. Despite some progress [Joseph et al.(2021)Joseph, Khan, Khan, and Balasubramanian, Gupta et al.(2022)Gupta, Narayan, Joseph, Khan, Khan, and Shah, Yang et al.(2022)Yang, Sun, Jiang, Xia, Zhang, Yuan, Wang, Luo, and Xu, Wu et al.(2022)Wu, Lu, Chen, Wu, Kang, and Yu, Ma et al.(2023c)Ma, Li, Zhang, Guo, Zhang, Gong, and Liu, Zohar et al.(2023b)Zohar, Wang, and Yeung, Ma et al.(2023a)Ma, Wang, Wei, Fan, Li, Liu, and Lv, Wang et al.(2023b)Wang, Yue, Hua, and Zhang, Sun et al.(2024)Sun, Li, and Mu] in this area, several key issues still remain unresolved. Many existing methods perform poorly at discovering unknown objects, leading to low recall for unknown classes. Additionally, existing OWOD methods [Joseph et al.(2021)Joseph, Khan, Khan, and Balasubramanian, Gupta et al.(2022)Gupta, Narayan, Joseph, Khan, Khan, and Shah, Sun et al.(2024)Sun, Li, and Mu, Wang et al.(2023b)Wang, Yue, Hua, and Zhang, Zohar et al.(2023b)Zohar, Wang, and Yeung, Ma et al.(2023c)Ma, Li, Zhang, Guo, Zhang, Gong, and Liu, Yang et al.(2022)Yang, Sun, Jiang, Xia, Zhang, Yuan, Wang, Luo, and Xu, Wu et al.(2022)Wu, Lu, Chen, Wu, Kang, and Yu, Ma et al.(2023a)Ma, Wang, Wei, Fan, Li, Liu, and Lv] employ a replay strategy during incremental learning, where data from previous tasks are reintroduced during the training of new classes. This strategy is suboptimal, leading to unnecessary consumption of computing resources and storage.
@@
Recently, the rise of visual-language pre-training has spawned a new area: open vocabulary object detection (OVD). In principle, OVD can detect novel classes defined by an unbounded (open) vocabulary at inference. Naturally, it is able to incrementally learn new classes by adding new prompts. With a limited vocabulary the model tends to misclassify near-out-of-distribution (NOOD) objects that have similar semantics to known classes, and ignore far-out-of-distribution (FOOD) objects. In real-world applications such as autonomous driving, where countless object-types cannot be initially included in the text prompt, OVD will inevitably fail when presented with OOD objects.
@@
To detect NOOD objects, we propose Multi-Scale Contrastive Anchor Learning (MSCAL). Contrastive learning has been used in OOD detection [Tack et al.(2020)Tack, Mo, Jeong, and Shin, Ming et al.(2023)Ming, Sun, Dia, and Li, Lu et al.(2024)Lu, Gong, Wang, Xue, Yao, and Moore] to pull similar samples together and push dissimilar samples apart in the representation space. The key insight of MSCAL is that in an open world setting, object classes are introduced gradually. As new classes are introduced, the decision boundaries for known classes shift within the shared feature space. We formulate the task of unknown object identification in the open-set setting as a series of deep one-class classification [Ruff et al.(2018)Ruff, Vandermeulen, Goernitz, Deecke, Siddiqui, Binder, Müller, and Kloft] problems. For each class $i$, we use an individual non-linear projector to map the feature pyramid into a class-specific representation space, contrasting embeddings with class anchors. Positive samples from class $i$, at different scales, maximize their similarity to the class anchor, while embeddings from other classes and the background act as negative samples. MSCAL ensures that embeddings of known classes at different scales are tightly clustered around their corresponding anchors, while unknown object embeddings are left out of the clusters and can be rejected based on distance.
Our method significantly surpasses state-of-the-art (SOTA) performance in U-Recall on the M-OWODB [Joseph et al.(2021)Joseph, Khan, Khan, and Balasubramanian] and S-OWODB [Gupta et al.(2022)Gupta, Narayan, Joseph, Khan, Khan, and Shah] benchmarks, while maintaining leading performance in other metrics. We further evaluate our method on a novel benchmark based on nuScenes [Caesar et al.(2019)Caesar, Bankiti, Lang, Vora, Liong, Xu, Krishnan, Pan, Baldan, and Beijbom], where it also achieves the best results. More importantly, our method preserves the zero-shot capability of the OVD model, as our implementation only optimizes text embeddings and additional MSCAL modules, while keeping the OVD model’s weights frozen. Therefore, our method provides a unified framework for both open vocabulary learning and open world learning. The contributions of this work are as follows:
@@ 2 Related Works
Closed set object detection has been extensively studied over the past decade [Girshick et al.(2014)Girshick, Donahue, Darrell, and Malik, Ren et al.(2015)Ren, He, Girshick, and Sun, Girshick(2015), Redmon et al.(2016)Redmon, Divvala, Girshick, and Farhadi, Liu et al.(2016)Liu, Anguelov, Erhan, Szegedy, Reed, Fu, and Berg, Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollar, Tian et al.(2019)Tian, Shen, Chen, and He, Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko]. To handle unknown objects and incrementally learn new classes, $\bmvaOneDot [Joseph et al$.(2021)Joseph, Khan, Khan, and Balasubramanian] first proposed an open world object detector, which extends the Faster R-CNN [Ren et al.(2015)Ren, He, Girshick, and Sun] with contrastive clustering and energy-based unknown-object identification. Subsequent works [Gupta et al.(2022)Gupta, Narayan, Joseph, Khan, Khan, and Shah, Ma et al.(2023c)Ma, Li, Zhang, Guo, Zhang, Gong, and Liu] used contextual information to improve unknown identification and knowledge transfer between known and unknown classes. To effectively detect unknown objects, $\bmvaOneDot [Zohar et al$.(2023b)Zohar, Wang, and Yeung] proposed Probabilistic Objectness (PROB) to estimate the objectness of different proposals. $\bmvaOneDot[Wang et al$.(2023b)Wang, Yue, Hua, and Zhang] introduced random proposals in detector training to encourage the unknown discovery and reduce confusion between known and unknown classes. $\bmvaOneDot[Ma et al$.(2023a)Ma, Wang, Wei, Fan, Li, Liu, and Lv] proposed decoupling object localization and classification via cascade decoding. Hyp-OW [Doan et al.(2024)Doan, Li, Behpour, He, Gou, and Ren] uses hyperbolic distance to enhance open world learning. $\bmvaOneDot[Sun et al$.(2024)Sun, Li, and Mu] further de-correlates objectness and class information by enforcing orthogonality. In some recent works [Ma et al.(2023b)Ma, Wang, Wei, Fan, Zhang, Sun, and Chen, Maaz et al.(2022)Maaz, Rasheed, Khan, Khan, Anwer, and Yang, Zohar et al.(2023a)Zohar, Lozano, Goel, Yeung, and Wang], foundation models are employed in open-world learning.
@@
While open vocabulary object detection (OVD) aims to detect novel classes with the help of vocabulary knowledge [Wu et al.(2024)Wu, Li, Xu, Yuan, Ding, Yang, Li, Zhang, Tong, Jiang, Ghanem, and Tao]. OVD models [Gu et al.(2022)Gu, Lin, Kuo, and Cui, Mal et al.(2022)Mal, Luo, Gao, Li, Chen, Wang, Zhang, and Hu, Rasheed et al.(2022)Rasheed, Maaz, Khattak, Khan, and Khan, Li* et al.(2022)Li*, Zhang*, Zhang*, Yang, Li, Zhong, Wang, Yuan, Zhang, Hwang, Chang, and Gao, Zhang et al.(2022)Zhang, Zhang, Hu, Chen, Li, Dai, Wang, Yuan, Hwang, and Gao, Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan, Wu et al.(2023)Wu, Zhu, Zhao, and Li, Du et al.(2022)Du, Wei, Zhang, Shi, Gao, and Li] typically match the image embeddings with the text embeddings. When presented with an object that is not in the prompt, the model will either assign an incorrect label of the nearest match (NOOD case) or no detection at all (FOOD case).
@@ 3 Method
Figure 1: Overview of the proposed method. During model training, we first initialize known class embeddings with a pretrained CLIP text encoder [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever]. The image encoder extracts a multi-scale feature map from the input. Then the RepVL-PAN [Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan] uses multi-level cross-modal fusion to combine image and text features, forming the feature pyramids. The detection head predicts the class label based on image-text similarity and regresses the bounding box. The detection loss is used to update the known class embeddings. Concurrently, MSCAL modules are trained to maximize the similarity between class anchor and spatial locations at different scales, and output a multi-scale score map to indicate whether an embedding is out-of-distribution (OOD) relative to a specified class. During the inference, the OOD map extracted by MSCAL is used to reduce known-unknown confusion. In addition, the pseudo unknown embedding used to discover unknown classes is constructed from the optimized known class embeddings and the generic “objectness” semantic concept.
@@
Open World Object Detection (OWOD) [Joseph et al.(2021)Joseph, Khan, Khan, and Balasubramanian] aims to detect both known and unknown objects while continuously learning new classes. At an arbitrary stage $t$, we consider the known classes as $\mathcal{K}^{t}=\{1,...,N\}$, and unknown classes as $\mathcal{U}$. An OWOD model should be able to detect objects in $\mathcal{K}^{t}$ and $\mathcal{U}$, and extend known classes to $\mathcal{K}^{t+1}=\mathcal{K}^{t}\cup\{N+1,...,N+k\}$ when $k$ new classes are incrementally learned. In this way, the object detector continuously discovers and learns new classes in the open world.
@@
Fig.˜1 shows the general architecture of the proposed method. Following [Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan], we use text $T$ and image $I$ as inputs and match the text embeddings with image embeddings to predict class labels and bounding boxes of objects. Let $W_{\mathcal{K}}=\{w_{1},...,w_{N}\}$ denote the text embedding of $N$ known classes, which is initialized from class names encoded by the pre-trained CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever] text encoder. $W_{\mathcal{K}}$ can be parameterized as model’s weight and optimized via Open World Embedding Learning (OWEL). During inference, a pseudo unknown embedding is constructed and appended to $W_{\mathcal{K}}$, and the CLIP text encoder is disposable. The image encoder (DarkNet backbone inherited from Yolo v8 [Jocher et al.(2023)Jocher, Chaurasia, and Qiu, Redmon and Farhadi(2018)]) extracts multi-scale features $C$ from the input image $I$. Then the multi-modal neck (RepVL-PAN [Cheng et al.(2024)Cheng, Song, Ge, Liu, Wang, and Shan]) uses multi-level cross-modal fusion to combine image and text features, forming the feature pyramids $P$. The detection head predicts bounding boxes and class labels by matching the cosine similarity of text embedding with each spatial location in $P$. Concurrently, MSCAL modules make dense predictions of out-of-distribution (OOD) scores and reject unknown bounding boxes in the detection head. Finally, the redundant predictions are filtered by Non-Maximum Suppression (NMS).
@@
Figure 2: Inferring the Pseudo Unknown Embedding in the embedding space. For CLIP-like models, text embeddings are mapped on a unit hypersphere.
Figure 3: MSCAL module. For each layer in the feature pyramid, all spatial locations will be mapped to a new space and contrasted with class anchors. The design of the projector follows [Wang et al.(2021)Wang, Zhou, Yu, Dai, Konukoglu, and Van Gool]. During inference, their inner product with the class anchor serves as the OOD score.
@@
Large pre-trained vision-language models have enabled many prompt learning methods [Zhou et al.(2022b)Zhou, Yang, Loy, and Liu, Zhou et al.(2022a)Zhou, Yang, Loy, and Liu, Agnolucci et al.(2023)Agnolucci, Baldrati, Todino, Becattini, Bertini, and Bimbo], which optimize text prompts rather than fine-tune the entire model to improve performance in downstream tasks. But they are not suitable for OWOD, because known $\mathcal{K}$ and unknown $\mathcal{U}$ classes are ever-changing. To address this, we propose a simple and effective way to learn new classes and detect unknown objects, called Open World Embedding Learning (OWEL). For $N$ known classes, we initialize known class embeddings $W_{\mathcal{K}}=\{w_{1},...,w_{N}\}$, and optimize them with the object detection loss. When $k$ new classes are introduced, we freeze $W_{\mathcal{K}}$ and train new class embeddings.
@@
When an OVD model is given an unknown object, it will either misclassify it as a semantically similar text prompt, or ignore it if the semantic difference is large. To avoid misclassification, we can model the distribution of known classes to reject out-of-distribution samples. To discover novel objects, we construct a Pseudo Class Embedding $w_{\mathcal{U}}$ representing unknown classes.
@@
Vision-language pretraining aligns the visual embedding space and the textual embedding space, such that the text embedding is equivalent to the corresponding image embedding [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever, Zhou et al.(2022b)Zhou, Yang, Loy, and Liu, Wang et al.(2023a)Wang, Li, Yao, and Li, Gan et al.(2022)Gan, Li, Li, Wang, Liu, and Gao]. It has been shown that the relation between text embeddings can be derived through vector offsets in continuous space language models [Mikolov et al.(2013b)Mikolov, Yih, and Zweig, Mikolov et al.(2013a)Mikolov, Chen, Corrado, and Dean]. As shown in Fig.˜3, let $\overline{w}$ denote the mean text embedding of all known classes, defined as:
[…]
@@
Let $w_{0}$ denote the text embedding representing generic objectness. Naturally, the text prompt of $w_{0}$ should be some general words, such as “object”, which is supported by a previous work [Maaz et al.(2022)Maaz, Rasheed, Khan, Khan, Anwer, and Yang] and our observation on YOLO-World. The semantics of $w_{0}$ overlaps with known classes to some extent. To shift its focus to unknown classes rather than making duplicate detections on known classes, we propose to construct a Pseudo Unknown Embedding, which is specialized to detect far-out-of-distribution (FOOD) objects. The Pseudo Unknown Embedding $w_{\mathcal{U}}$ is defined by subtracting the mean of known classes from $w_{0}$:
$w¯,w_{\mathcal{U}}=w_{0}-\alpha\frac$$\overline$$|\overline{w}$||}, (2)